OpenStudy (anonymous):

Add the fractions. Select the answer that is in simplest form. 2/5 + 3/10 =?

OpenStudy (anonymous):

7/10

OpenStudy (anonymous):

2/5= 4/10

OpenStudy (anonymous):

4+3= 7

OpenStudy (anonymous):

where did the 4 come from

OpenStudy (anonymous):

2/5 multi both 2 and 5 by 2

OpenStudy (anonymous):

2/5 = 4/10